# D2USt3R: Enhancing 3D Reconstruction with 4D Pointmaps for Dynamic Scenes

### v1

@@ B. Bullet time reconstruction for dynamic video input
Owing to dynamic alignment and 4D pointmaps, we can aggregate and render a highly dynamic video input into a single, coherent bullet-time view. As illustrated in Figure 10, a video consisting of large input frames can be aligned into an intermediate bullet-time representation. This approach enables static reconstruction even when the input video includes dynamic motion. Consequently, it becomes feasible to directly train a 3D Gaussian splatting [20] on scenes containing moving people in landmarks or dynamics that are challenging to render using conventional methods.
@@ C. Additional results
In Figure 11 and Figure 12, we additionally present qualitative results for pointmap reconstruction and depth estimation. Utilizing our 4D pointmaps, dynamic elements are well-aligned, thereby enhancing both depth estimation and 3D reconstruction quality. In Figure 13 and Figure 14, we present qualitative results for optical flow on DAVIS [30] and KITTI [11] datasets. We observed that D2USt3R is capable of accurately predicting optical flow without relying on any dedicated optical flow module.

### v2

@@ B. Bullet time reconstruction for dynamic video input
Owing to dynamic alignment and SDAP, we can aggregate and render a highly dynamic video input into a single, coherent bullet-time view. As illustrated in Figure 10, a video consisting of large input frames can be aligned into an intermediate bullet-time representation. This approach enables static reconstruction even when the input video includes dynamic motion. Consequently, it becomes feasible to directly train a 3D Gaussian splatting kerbl20233d on scenes containing moving people in landmarks or dynamics that are challenging to render using conventional methods.
@@ C. Additional results
In Figure 11 and Figure 12, we additionally present qualitative results for pointmap reconstruction and depth estimation. Utilizing our SDAP, dynamic elements are well-aligned, thereby enhancing both depth estimation and 3D reconstruction quality. In Figure 13 and Figure 14, we present qualitative results for optical flow on DAVIS pont20172017 and KITTI geiger2013vision datasets. We observed that D2USt3R is capable of accurately predicting optical flow without relying on any dedicated optical flow module.
@@
D. End-to-end dynamic mask head integration
Table 9: Multi-frame depth estimation results on end-to-end integration analysis.
Methods TUM-Dynamics Bonn Sintel KITTI
[…]
D2USt3R 0.142 83.9 0.060 95.8 0.324 57.5 0.104 90.7
D2USt3R + Dynamic mask head 0.140 84.4 0.059 95.6 0.313 58.3 0.105 90.3
We further explored an end-to-end training strategy in which the entire network, including the dynamic mask head, is jointly optimized. As shown in Table 9, this integrated setup encourages implicit interactions between sub-tasks through shared representations, leading to consistent, albeit moderate, performance improvements. Specifically, we observe gains on two of the four benchmarks, with a mean reduction in AbsRel of 1.7% and a mean increase in $\delta$ of 0.9 percentage points. These results confirm that joint optimization promotes beneficial cross-task interactions, yielding complementary improvements across depth and dynamic mask estimation.
E. Limitations
Despite effectively handling dynamic scenarios, our dynamic alignment process has certain limitations. Since our model relies on an off-the-shelf optical flow model, its performance is dependent on the quality of the predicted optical flow. Additionally, our model is non-generative and can only take two frames as input. Because we aim to achieve alignment across time, extending our approach to video input would require an additional global optimization technique.